Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4-mph current, it took her 20 minutes longer to get there than to return. How fast will her boat go in still water?
Answer by stanbon(75887)

Debbie traveled by boat 5 miles upstream to fish in her favorite spot. Because of the 4-mph current, it took her 20 minutes longer to get there than to return. How fast will her boat go in still water?
------------------------------------------------------
Upstream DATA:
distance = 5 miles; rate = b-4 mph ; time = d/r = 5/(b-4) hrs
----------------------------------------------------------------------
Downstream DATA:
distance = 5 miles; rate = b+4 mph ; time = 5/(b+4) hrs.
----------------------------------------------------------------------
Equation:
uptime - downtime = 1/3 hr
5/(b-4) - 5/(b+4) = 1/3
5(b+4) - 5(b-4) = (1/3)(b^2-16)
40 = (1/3)(b^2-16)
b^2 - 16 = 120
b^2 = 136
---
Positive solution:
b = 11.66 mph (boat speed in quiet water)
==============================
Cheers,
Stan H.